OpenStudy (chupacabraj):

How do you graph these equations in 3D?

OpenStudy (chupacabraj):

\[z_1=x^2+y^2\] and \[z_2=2-x^2-y^2\]

OpenStudy (chupacabraj):

@zepdrix

zepdrix (zepdrix):